OpenStudy (anonymous):

The formula A = 118e^(0.024t) models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? A. 2008 B. 2005 C. 2006 D. 2007 Im not sure how to figure out t..

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = 118e^(0.024t) 140 = 118e^(0.024t) Now solve for t

OpenStudy (anonymous):

The formula A = 118e^0.024t models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? 140 = 118 e^0.024t 140/118 = e^0.024t ln(140/118) = 0.024t t = ln(140/118)/0.024 = 7 years 1998 + 7 = 2005 B. 2005

OpenStudy (anonymous):

ohh okay that makes so much sense! thanks!

OpenStudy (anonymous):

Ur welcome